• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Shabestari

    Log entry time 10:45:16 on December 03, 2007

    Entry number 218192

    keyword=Shift summary, Tuesday 11/27/2007 (Problem with HALOG)

    Shift summary (Tuesday 11/27/2007)

    16:00 No beam, target has moved to empty 17:15 He target vented, got controlled access to hall 18:00 J.P. still in the hall working on the target 18:16 He target filled and we are in beam permit. No beam yet. 18:35 MCC called. Tuning the beam. 20:55 No beam 21:14 MCC called. They are ready to send CW, and run the harp scans 22:45 MCC called. Harp scans result: A: 217 x 208 micron B: 200 x 100 micron 23:20 raster scan looks fine (2- mm x 2+ mm) 23:25 we are still unable to make log entry. For harp scans refer to elog entry # 1403220 23:27 MCC called. They are taking off the raster for a while and will put it back shortly. 23:45 Beam is back. Moving target to right Iron



    A copy of this log entry has been emailed to: mitras@jlab.org, vasulk@jlab.org, jpchen@jlab.org